# A Graph-Theoretical Framework to Analyse Zero Discord Quantum States

### v1

@@ 3 Quantum Discord of weighted Graphs
Proof: Let $(G,a)$ be a weighted graph defined on ${\mathbb{C}}^{2}\otimes{\mathbb{C}}^{2}$. ${\mathbb{C}}^{2}\otimes{\mathbb{C}}^{2}$ and $\rho_{G}=[\rho_{ij}]_{4\times 4}$ represents the density operator. Clearly, $(\overline{\text{Partial}U)\rho_{G}}=\rho_{G}$ implies that blocks in the density operator are Hermitian matrices. Moreover, $(\rho_{11}-\rho_{22})\rho_{14}=\rho_{12}(\rho_{13}-\rho_{24})$ reveals commutativity of the blocks as expressed by $\rho_{11}\rho_{14}+\rho_{12}\rho_{24}=\rho_{12}\rho_{13}+\rho_{22}\rho_{14}$. Hence the proof [16].
@@
The quantum discord $D_{Q}(G,a)=0$ if $\sum_{k}{A_{ik}}^{xy}{A_{kj}}^{rs}=\sum_{k}{{A_{ik}}^{rs}{A_{kj}}^{xy}}$ for $i,j$ and $\overline{Partial(U^{i})\rho_{G}}=\rho_{G}$, for $U^{1}=\underbrace{I_{2}\otimes\dots\otimes I_{2}}_{p}\otimes U_{p+1}\dots\otimes U_{p+q}$, and $U^{2}=I_{2}\otimes U_{2}\dots\otimes U_{n}$, where, $U_{k}$ are either $I_{2}$ or $\sigma_{x}$.
Proof: Consider a weighted graph $(G,a)$ with $2^{n}$ vertices related to an $n$-qubit quantum state. Let $\rho_{G}=[A^{xy}]{2\times 2}$ represent the density operator of this graph, where $A^{xy}$ is a block matrix of size $2^{n-q-1}$ and $A^{xy}$ is composed of elements denoted as $A{ij}^{xy}$. If $\sum_{k}{A_{ik}}^{xy}{A_{kj}}^{rs}=\sum_{k}{{A_{ik}}^{rs}{A_{kj}}^{xy}}$ for $i,j$ and $\overline{Partial(U^{i})\rho_{G}}=\rho_{G}$, for $U^{1}=\underbrace{I_{2}\otimes\dots\otimes I_{2}}_{p}\otimes U_{p+1}\dots\otimes U_{p+q}$, and $U^{2}=I_{2}\otimes U_{2}\dots\otimes U_{n}$, where, $U_{i}$ are either $I_{2}$ or $\sigma_{x}$, then $A^{xy}$ and $A^{mn}$ commute, and each block of $A^{xy}$ also commutes. Hence, the discord is zero [16].

### v2

@@ 3 Quantum Discord of weighted Graphs
Proof: Let $(G,a)$ be a weighted graph defined on ${\mathbb{C}}^{2}\otimes{\mathbb{C}}^{2}$ and $\rho_{G}=[\rho_{ij}]_{4\times 4}$ represents the density operator. Clearly, $(\overline{\text{Partial}U)\rho_{G}}=\rho_{G}$ implies that blocks in the density operator are Hermitian matrices. Moreover, $(\rho_{11}-\rho_{22})\rho_{14}=\rho_{12}(\rho_{13}-\rho_{24})$ reveals commutativity of the blocks as expressed by $\rho_{11}\rho_{14}+\rho_{12}\rho_{24}=\rho_{12}\rho_{13}+\rho_{22}\rho_{14}$. Hence the proof [16].
@@
The quantum discord $D_{Q}(G,a)=0$ if $\sum_{k}{A_{ik}}^{xy}{A_{kj}}^{rs}=\sum_{k}{{A_{ik}}^{rs}{A_{kj}}^{xy}}$ for all $i,j$ and $\overline{Partial(U^{i})\rho_{G}}=\rho_{G}$, for $U^{1}=\underbrace{I_{2}\otimes\dots\otimes I_{2}}_{p}\otimes U_{p+1}\dots\otimes U_{p+q}$, and $U^{2}=I_{2}\otimes U_{2}\dots\otimes U_{n}$, where, $U_{k}$ are either $I_{2}$ or $\sigma_{x}$.
Proof: Consider a weighted graph $(G,a)$ with $2^{n}$ vertices related to an $n$-qubit quantum state. Let $\rho_{G}=[A^{xy}]{2\times 2}$ represent the density operator of this graph, where $A^{xy}$ is a block matrix of size $2^{n-q-1}$ and $A^{xy}$ is composed of elements denoted as $A{ij}^{xy}$. If $\sum_{k}{A_{ik}}^{xy}{A_{kj}}^{rs}=\sum_{k}{{A_{ik}}^{rs}{A_{kj}}^{xy}}$ for all $i,j$ and $\overline{Partial(U^{i})\rho_{G}}=\rho_{G}$, for $U^{1}=\underbrace{I_{2}\otimes\dots\otimes I_{2}}_{p}\otimes U_{p+1}\dots\otimes U_{p+q}$, and $U^{2}=I_{2}\otimes U_{2}\dots\otimes U_{n}$, where, $U_{i}$ are either $I_{2}$ or $\sigma_{x}$, then $A^{xy}$ and $A^{mn}$ commute, and each block of $A^{xy}$ also commutes. Hence, the discord is zero [16].